OpenStudy (anonymous):

by using the discriminant, determine the value for k such that the equation will have exactly one root.

OpenStudy (anonymous):

\[2x ^{2}-kx+3=0\]

OpenStudy (amistre64):

when k = 0 id assume :)

OpenStudy (amistre64):

wrong "k" lol sorry

OpenStudy (anonymous):

you want the discriminant equal to 0

OpenStudy (amistre64):

k^2 -4ac = 0 k^2 -4(2)(3) = k^2 - 24

OpenStudy (anonymous):

ha I supose. I know the answer is k= +/-2 (root6) I just dont know how they got it.

OpenStudy (amistre64):

k = +- sqrt(24) k = +- sqrt(4*6) k = +- 2sqrt(6)

OpenStudy (anonymous):

ohh I see! THanks